Mal87:

y varies inversely with x. If x is 6.4 when y is 1.5, what is k, the constant of inverse variation? Round your answer to the nearest hundredth, if necessary.

Mal87:

This is what I have so far y=kx k=y*x=1.5*6.4=9.60\]

Hero:

@Mal87, you posted the formula for direct variation. The problem posted indicates that situation is one where y has an inverse relationship with x rather than a direct one.